ONE GRAM OF GOLD CAN BE BEATEN OUT INTO A FOIL 1.O M2 IN AREA. HOW MANY ATOMS THICK IS SUCH A FOIL? THE MASS OF GOLD ATOM IS 3.27 x 10-25 KG , DENSITY OF GOLD 1.9 X 104 KG.


1
Expert's answer
2021-02-18T18:48:13-0500

Given quantities:

"m = 1g \space\space\space\space\space S = 1m^2 \space\space\space\space \rho = 19000 \large\frac{kg}{m^3}" "m_o=" 3.27 x 10-25 KG

"N = \large\frac{m}{m_o}" "= \frac{1*10^{-3}}{3.27*10^{-25}}=3*10^{21}"